If G one earth is 3kN. How much is m and G on the moon if the gravity there is 1.62m/s²

1
Expert's answer
2021-12-22T14:10:07-0500

The mass on the earth is

m=G/g=3000/9.8=306 kg.


The mass on the moon remains the same because this is an intrinsic property of the object, while the weight will be

GM=gM•m=1.62•306=496 N.